What is the x intercept of 7x-8y=0

Answer by nyc_function(2741)

7x - 8y = 0

To find the x-intercept, set y = 0 and solve for x.

7x - 8(0)= 0

7x - 0 = 0

7x = 0

x = 0/7

x = 0